# Anticipating Gaming to Incentivize Improvement: Guiding Agents in (Fair) Strategic Classification

### v1

@@ A.5 Type 3 Experiments
We next consider a population consisting of two equal-size groups a and b, with group a being majority-qualified ($\alpha_{a}$ = 0.7) and group b being majority-unqualified ($\alpha_{b}$ = 0.2). The feature distributions for both groups follow truncated Gaussian distributions, with $G^{0}_{a}\sim\text{TruncNorm}([20,110],65,15^{2}),G^{0}_{b}\sim\text{TruncNorm}([0,90],45,15^{2}),G^{1}_{a}\sim\text{TruncNorm}([98,188],143,15^{2}),$ and $G^{1}_{b}\sim\text{TruncNorm}([78,168],123,15^{2})$. We assume both groups have access to the same strategic actions, with $C_{M,s}=0.2$, $C_{I,s}=0.8$, and the same boost distribution as in the "Single-group experiment".
@@
Unlike Type 1, comparing DP vs. TPR fairness impacts reveals greater threshold changes under DP for both firms, as seen in Fig. 11. Without improvement actions, $\alpha_{s}$ remains unchanged, forcing firms to lower $\theta_{b}$ and raise $\theta_{a}$ to meet DP constraints, as strategic responses can’t achieve selection parity. For group $b$, the DP-fair strategic firm increases unqualified manipulation more, raising $\bm{\Psi}^{0}_{b,(1)}(\hat{\theta}^{\mathcal{C}}_{b})$ compared to the unfair policy’s high threshold (curb many unqualified manipulation), shown by solid red lines in Figure 12’s first plot. The non-strategic firm lowers $\theta_{b}$ similarly, but with less manipulation increase (Fig. 12’s second plot) due to already high unqualified manipulation. Thus, without improvement actions, fairness requires larger threshold shifts under strategic responses, reducing firm utility.

### v2

@@ D.2 Synthetic dataset Type 3 Experiments
We next consider a population consisting of two equal-size groups a and b, with group a being majority-qualified ($\alpha_{a}$ = 0.7) and group b being majority-unqualified ($\alpha_{b}$ = 0.2). The feature distributions for both groups follow truncated Gaussian distributions, with $G^{0}_{a}\sim\text{TruncNorm}([20,110],65,15^{2}),G^{0}_{b}\sim\text{TruncNorm}([0,90],45,15^{2}),G^{1}_{a}\sim\text{TruncNorm}([98,188],143,15^{2})$ and $G^{1}_{b}\sim\text{TruncNorm}([78,168],123,15^{2})$. We assume both groups have access to the same strategic actions, with $C_{M,s}=0.2$, $C_{I,s}=0.8$, and the same boost distribution as in the "Single-group experiment".
@@
Unlike Type 1, comparing DP vs. TPR fairness impacts reveals greater threshold changes under DP for both firms, as seen in Figure 22. Without improvement actions, $\alpha_{s}$ remains unchanged, forcing firms to lower $\theta_{b}$ and raise $\theta_{a}$ to meet DP constraints, as strategic responses can’t achieve selection parity. For group $b$, the DP-fair strategic firm increases unqualified manipulation more, raising $\bm{\Psi}^{0}_{b,(1)}(\hat{\theta}^{\mathcal{C}}_{b})$ compared to the unfair policy’s high threshold (curb many unqualified manipulation), shown by solid red lines in Figure 23’s first plot. The non-strategic firm lowers $\theta_{b}$ similarly, but with less manipulation increase (Figure 23’s second plot) due to already high unqualified manipulation. Thus, without improvement actions, fairness requires larger threshold shifts under strategic responses, reducing firm utility.